OpenStudy (anonymous):

What is the geometric mean of 7 and 12? 1 9.5 2√21 4√21<--

OpenStudy (anonymous):

Checking answers and the first choice is 1 5/7

Parth (parthkohli):

hmm, how did you arrive at that answer?

OpenStudy (anonymous):

7 x 12 = 84 square root of 84 should be 4 square root 21

Parth (parthkohli):

why?

OpenStudy (anonymous):

4 x 21 = 84 and then you simplify it further so the answer is actually 2 square root 21

Parth (parthkohli):

yes

OpenStudy (anonymous):

Cool thanks

OpenStudy (anonymous):

Let G be the geometric mean of 7 and 12 \[G^2 = 7 \times 12 \rightarrow G^2 = 84 \rightarrow G = \sqrt{84 }\] \[\rightarrow G = \sqrt{84 } \rightarrow G = \sqrt{4 \times 21} \rightarrow G = 2 \sqrt{21} \] @ILoVeMaTh3